# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ 5.7 BFFs and high-energy neutrinos
AT2019aalc has been associated with the high-energy neutrino event IC-191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al. 2024). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1; Abbasi et al. 2023). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC-230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Figs. 24 and 24. We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
@@
The neutrino associated with AT2021loi is days delayed with respect to its first optical peak and days with respect to its second optical peak. Note, that the nuclear transients (including AT2019aalc) are associated with neutrinos that were delayed by days, which might be explained with delayed mass accretion, delayed jet formation or outflow ejection (van Velzen et al. 2024). In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al. 2023) is compatible with a newly ejected jet or outflow component, similarly to the case of AT2019aalc. However, from the current data, we cannot tell if the radio-emitting region of AT2019aalc has been ejected before or after the neutrino. Although we do not see a clear increase in the MHz RACS data months after the neutrino arrival. However, if the radio flare is indeed related to a newly ejected, synchrotron self-absorbed component, it is not surprising that it does not result in any increasing at low frequencies at early times. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al. 2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients. These findings indicate that outflows play an important role in the neutrino production. Moreover, Winter & Lunardini (2023) estimate the highest neutrino fluence for AT2019aalc among the studied nuclear transients associated with neutrinos. A similar case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
@@
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al. 2024; Hinkle 2024; see also Subsect. 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al. 2023; see also Subsect. 5.1).
@@ 6 Conclusions
We conducted a multi-wavelength campaign during the second optical flare of the nuclear transient AT2019aalc associated with a high-energy neutrino event. We found that the second optical flare is even more luminous and decays even slower than the first one. Both flares evolve rapidly and with constant color. The optical flare is accompanied by UV emission which peaks at extremely high luminosities (at erg/s). The IR emission increases, similar to the first flare when it peaked delayed with respect to the optical and was explained with a dust echo flare. We detected a very soft X-ray flare that peaked before the optical/UV emission peaks. Later, the Swift/XRT detected an almost times more luminous X-ray flare which is clearly dominated by a blackbody component, suggesting enhanced accretion or disk instability.
Our VLBI observation provides a hint of extension of the radio-emitting source and clearly favors an AGN-like origin over a star-formation scenario. The radio monitoring revealed a long-term flare and an inverted spectrum above GHz, which is most probably explained by a newly ejected, self synchrotron-absorbed outflow or jet component. As we do not see obvious signs of variability well before the first optical flare, we suggest that the transient event in renewed the activity of the host AGN. Restarted AGN tend to have inverted radio spectrum. This also naturally explains the long-term flare.
The results from optical spectroscopy help us to better understand the event responsible for restarting the AGN. In the optical spectra, the presence of persistent Bowen lines are consistent with intensified accretion over a year after the second flare. The UV bright nature of the transient is consistent with the Bowen mechanism which pumps lines mostly in this range. The line intensities of the Bowen lines and the temperature of the BLR both peak around the peak of the optical flare indicating that the optical flare is driven by the ionization of the clouds of most probably the BLR. As we find the X-ray flaring episodes prior to the rise in optical/UV, the origin of the latter is most probably reprocessed X-ray photons which are also required for the extreme ionization implied by the Bowen lines. The X-ray flares are very soft which explains the unusually strong He ii lines. These findings altogether suggest the accretion-driven origin of the flaring episode we monitored. Note, that reprocessed soft X-ray emission i.e. by an outflow surrounding the disk is a possible explanation for the origin of the optical/UV emission of TDEs (Bu et al. 2022).
We detected several persistent or even late-time increasing high-ionization coronal lines in the spectra of AT2019aalc. The detection of [S xii requires an extremely high ionization potential. To our knowledge, AT2019aalc is only the third extragalactic transient event detected with this line and only Seyfert galaxies (including the neutrino emitter NGC 1068) and ECLEs were reported with a line with such a high ionization potential. The Fe/X-ray ratios disfavor a past TDE interpretation responsible for the coronal lines, which was proposed to explain these lines seen in most of the ECLEs studied earlier. The slowly decaying optical emission might instead be explained with continuous photoionization of the dust grains (located at distances consistent with the BLR instead of the NLR as seen for ECLEs before). This is consistent with the early observation of the dust echo emission with respect to the optical peak which suggests that the dusty region is located very close to the SMBH. Based on the narrow-line ratios, we found that enhanced star formation activity is a favorable explanation for the dusty environment. Moreover, the gas-rich environment might explain the restarted activity of the AGN.
The high [Fe x / [O iii luminosity ratios of the BFFs, coronal line-detected TDEs and ECLEs suggest a possible connection between these source types. These sources tend to have large dust echo-like flares, which implies that the liberation of Fe from the dust grains can be explained via their TDE(-like) flaring episodes. The significant photoionization is supported by the detection of strongly ionized neon lines in the EELR. A delayed outflow/jet component and collisional excitation provides an alternative explanation for the coronal lines and their late-time increasing or appearance. The BFF and coronal-line emitter AT2021loi has an inverted radio spectrum and based on survey data we found a flaring radio source at the position of AT2017bgt. These findings suggest that newly ejected radio-emitting components play an important role not only in BFFs, but might be connected to the coronal line emission. A detailed radio study of other BFFs and coronal line emitters is crucial to better understand the appearance of these highly ionized lines.
Altogether, our results are consistent with the Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter. A connection between the ECLEs, some TDEs and BFFs suggested by our spectroscopy results makes a non-standard TDE-related scenario more probable to explain the optical flares and i.e. the restarted activity of the host AGN. Moreover, due to the pre-existing BLR and therefore the high densities, TDE-AGN are more capable for the Bowen Flurosence mechanism and the high-ionization coronal lines. Only a few TDEs were detected with coronal lines; these are also known as TDEs with extreme dust echo emission.
Continuous interactions of a pair of binary star and the SMBH favored by Trakhtenbrot et al. (2019) might explain the re-brightened and slowly decaying optical emission of BFFs. An alternative scenario for the two distinct optical flares is a partial disruption of a single star in an AGN. Here, the slowly decaying reprocessed optical/UV emission originates from soft X-ray flares powered by continuous tidal interaction between the stellar debris and the pre-existing accretion disk. The double-peaked optical light curve is explained with a second disruption due to the surviving core after the first passage. However, more BFFs are needed to study the mechanism responsible for these transients. TDE-related scenarios nevertheless provide an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the ionized gas. Future observations will help us to better understand the nature of the transient; a binary TDE would not produce a third flare, a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
@@
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC-230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al. (2024). Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionozation coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode/bumps in optical. In both cases the host galaxies are classified as broad-line AGN with remarkable star formation activity.
Acknowledgements
We thank Benny Trakhtenbrot for useful discussions. PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal. EH acknowledges support by NASA under award number 80GSFC21M0002.
@@ Acknowledgements
The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: EV027. e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. We acknowledge the use of public data from the Swift data archive. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. The Australia Telescope Compact Array is part of the Australia Telescope National Facility https://ror.org/05qajvd42 which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the Observatory site. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. These results made use of the Lowell Discovery Telescope (LDT) at Lowell Observatory. Lowell is a private, non-profit institution dedicated to astrophysical research and public appreciation of astronomy and operates the LDT in partnership with Boston University, the University of Maryland, the University of Toledo, Northern Arizona University and Yale University. The upgrade of the DeVeny optical spectrograph has been funded by a generous grant from John and Ginger Giovale and by a grant from the Mt. Cuba Astronomical Foundation. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.

### v2

@@ 5.8 BFFs and high-energy neutrinos
AT2019aalc was associated with the high-energy neutrino event IC191119A (most probable neutrino energy: TeV) detected by the IceCube Observatory (IceCube Collaboration 2019; van Velzen et al. 2024). Motivated by the classification of AT2019aalc suggested by its multi-wavelength properties (Subsect. 5.6), we searched for spatial coincidences between the earlier classified BFFs and the neutrino alerts from the IceCube Event Catalog of Alert Tracks (ICECAT-1: Abbasi et al. 2023). We found that the BFF AT2021loi is located within the 90% rectangular uncertainty contour of the high-energy neutrino event IC230511 (most probable neutrino energy: TeV) detected in May (IceCube Collaboration 2023). In the full likelihood contours published in ICECAT-1, the source is just outside the 90% contour. We present the source positions together with the neutrino best-fit positions and both contours in Fig 16. We also report a coincidence between the ’prototype’ BFF AT2017bgt and the high-energy neutrino event IC-200410A. This event is, however, a poorly reconstructed cascade event, therefore we do not discuss it further here.
@@
The neutrino associated with AT2021loi was detected days and days after the first and second optical peaks, respectively. Notably, all three nuclear transients (including AT2019aalc) associated with neutrinos exhibited a similar pattern, with neutrino detection occurring days after the optical peaks. This might be explained with delayed outflow activity (van Velzen et al. 2024). In the case of AT2021loi, the inverted radio spectrum (Makrygianni et al. 2023) is compatible with a newly ejected outflow component, similarly to the case of AT2019aalc. Interactions between the out-flowing material and clouds in the surrounding of the SMBH (Wu et al. 2022) or between ultra-high energy protons (accelerated in a jet or outflow) and infrared photons from the dust echo (Winter & Lunardini 2023) were proposed to explain the observed neutrino emission of the nuclear transients. These findings indicate that outflows play an important role in the neutrino production. A case study for AT2021loi would help us to further constrain the neutrino production models of nuclear transient sources.
@@
If the connection between the high-ionization coronal line detection and the unusually large dust echos of nuclear transients is real, this can be used to reveal more sources with strong dust echo and therefore more candidate neutrino sources. Notably, both AT2019aalc and AT2021loi are characterized by significant dust echos (van Velzen et al. 2024; Hinkle 2024: see also Subsect. 4.2) and unusually strong high-ionization coronal lines (Makrygianni et al. 2023: see also Subsect. 5.1).
@@ 6 Conclusions
We conducted a multi-wavelength campaign during the re-brightening of the nuclear transient AT2019aalc associated with a high-energy neutrino event. We found that the second optical flare is more luminous and decays slower than the first one. Both flares evolve rapidly and with constant color. The UV emission peaks at extremely high luminosities (at erg/s). The re-brightening is accompanied by an IR flare, similarly to the first flare when it was explained with dust echo emission. An X-ray flare peaked before the optical/UV emission, preceding two subsequent flares that were nearly an order of magnitude brighter and predominantly blackbody in nature. The X-ray flares are very soft which explains the unusually strong He ii lines. In the optical spectra, persistent Bowen fluorescence lines were detected over a year after the second flare. With VLBI we clearly detect a non-thermal radio source which favors an AGN-like origin. The radio flux monitoring revealed a long-term flare and an inverted spectrum above GHz, which is likely explained by a newly ejected, self synchrotron-absorbed outflow component. Altogether, our results suggest a Bowen Fluorescence Flare classification of AT2019aalc. We also classify AT2019aalc as an Extreme Coronal Line Emitter.
The two distinct optical flares might be attributed to the partial disruption of a single star within an AGN environment. The stellar debris-disk interaction in such a system could result in complex light curve variations e.g., the plateau observed in optical/UV (Chan et al. 2020). A TDE-related scenario provides an explanation not only for the suddenly enhanced activity of the nucleus but also for the high metallicity of the disk gas which was notably suggested by TDE-AGN simulations (Ryu et al. 2024). Future observations will help us to better understand the nature of AT2019aalc: a repeating TDE would lead to a roughly periodic flaring, while some other AGN activity would lead to erratic future flares.
@@
Motivated by the high-energy neutrino associated with AT2019aalc, we cross-matched the ICECAT-1 with the positions of the known 4 BFFs. We found that AT2021loi is coincident with the high-energy neutrino event IC230511 confirming the flaring AGN-neutrino connection studied by van Velzen et al. (2024). Apart from the known multi-wavelength properties of BFFs, AT2021loi shares characteristics with AT2019aalc such as late-time IR flare (due to dust echo), high-ionization coronal lines in its spectra (and also -band excess), radio detection (and inverted radio spectrum) and re-brightening episode in optical. In both cases the AGN are classified as broad-line AGN in-prior the transients occurred.
Acknowledgements.
@@
We thank Benny Trakhtenbrot for useful discussions. PMV, AF, BA, ST, AK, DJB, EM-B acknowledge the support from the DFG via the Collaborative Research Center SFB1491 Cosmic Interacting Matters - From Source to Signal. EH acknowledges support by NASA under award number 80GSFC21M0002.The European VLBI Network is a joint facility of independent European, African, Asian, and North American radio astronomy institutes. Scientific results from data presented in this publication are derived from the following EVN project code: EV027. e-MERLIN is a National Facility operated by the University of Manchester at Jodrell Bank Observatory on behalf of STFC. We acknowledge the use of public data from the Swift data archive. This research has made use of the XRT Data Analysis Software (XRTDAS) developed under the responsibility of the ASI Science Data Center (ASDC), Italy. The Australia Telescope Compact Array is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Gomeroi people as the Traditional Owners of the Observatory site. Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including Caltech, The Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, The Flatiron Institute, the Gotham Participation Group, Harvard University, Heidelberg University, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, and NEOWISE, which is a project of the Jet Propulsion Laboratory/California Institute of Technology. WISE and NEOWISE are funded by the National Aeronautics and Space Administration. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The Asteroid Terrestrial-impact Last Alert System (ATLAS) project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The Pan-STARRS1 Surveys (PS1) and the PS1 public science archive have been made possible through contributions by the Institute for Astronomy, the University of Hawaii, the Pan-STARRS Project Office, the Max-Planck Society and its participating institutes, the Max Planck Institute for Astronomy, Heidelberg and the Max Planck Institute for Extraterrestrial Physics, Garching, The Johns Hopkins University, Durham University, the University of Edinburgh, the Queen’s University Belfast, the Harvard-Smithsonian Center for Astrophysics, the Las Cumbres Observatory Global Telescope Network Incorporated, the National Central University of Taiwan, the Space Telescope Science Institute, the National Aeronautics and Space Administration under Grant No. NNX08AR22G issued through the Planetary Science Division of the NASA Science Mission Directorate, the National Science Foundation Grant No. AST-1238877, the University of Maryland, Eotvos Lorand University (ELTE), the Los Alamos National Laboratory, and the Gordon and Betty Moore Foundation. These results made use of the Lowell Discovery Telescope (LDT) at Lowell Observatory. Lowell is a private, non-profit institution dedicated to astrophysical research and public appreciation of astronomy and operates the LDT in partnership with Boston University, the University of Maryland, the University of Toledo, Northern Arizona University and Yale University. The upgrade of the DeVeny optical spectrograph has been funded by a generous grant from John and Ginger Giovale and by a grant from the Mt. Cuba Astronomical Foundation. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Native Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.
Appendix A Spetroscopic details
Date | Telescope | Instrument | Exp. time | Slit width | Seeing FWHM |
(s) | (arcsec) | (arcsec) | |||
Post-flare-1 | |||||
2021-07-06 | Keck-I | LRIS | 300 | 1.0 | 1.1 |
Over-flare-2 | |||||
2023-06-15 | Keck-I | LRIS | 300 | 1.0 | 1.4 |
2023-06-22 | LDT | DeVeny | 800 | 1.5 | 1.2 |
2023-07-11 | Palomar Hale m | DBSP | 180 | 1.5 | 1.5 |
2023-08-17 | Palomar Hale m | DBSP | 900 | 1.5 | 1.1 |
2024-05-21 | Palomar Hale m | DBSP | 1200 | 1.5 | 1.3 |
Date | Instrument | Resolving power | |||
(approx.) | (days) | ||||
2008-04-05 | SDSS | 1500-2500 | -11 years | 2a,b | |
2021-06-07 | LRIS | 1000 | -737 | 0.21 0.01 | 4a,b,c,d |
2023-06-15 | LRIS | 1000 | -28 | 0.64 0.02 | 4a,b,c,d |
2023-06-22 | Deveny | 750 | -21 | 0.50 0.01 | 4a,b,c,d |
2023-07-11 | DBSP | 1000 | -2 | 0.83 0.04 | 4a,b,c,d |
2023-08-17 | DBSP | 1000 | +35 | 0.25 0.01 | 5a,b,c,d,e |
2024-05-21 | DBSP | 1000 | +315 | 0.30 0.02 | 5a,b,c,d,e |